OpenStudy (anonymous):

is the sum of a series equal to what it converges to? Find the sum of the convergent series.

OpenStudy (anonymous):

\[\sum_{n=0}^{\infty}5/-2^n\]

OpenStudy (anonymous):

the answer is got was -infinity but the real answer is -5/3

OpenStudy (anonymous):

Do you mean \[\sum_{n=0}^\infty\left(-\frac{5}{2}\right)^n\quad\text{or}\quad\sum_{n=0}^\infty5\left(-\frac{1}{2}\right)^n\quad\text{or something else altogether?}\]

OpenStudy (anonymous):

\[\sum_{n=0}^{\infty}(\frac{ -5 }{ 2^{n}})\]

OpenStudy (anonymous):

So the exponent doesn't apply to the -5, correct?

OpenStudy (anonymous):

The exponent does not apply to -5.

OpenStudy (anonymous):

Alright, so \[\sum_{n=0}^\infty -\frac{5}{2^n}=-5\sum_{n=0}^\infty \left(\frac{1}{2}\right)^n\] which is a geometric series with common ratio \(\dfrac{1}{2}\). Since the magnitude of the common ratio is less than 1, you know the series converges. For a convergent geometric series of the form \[\sum_{n=0}^\infty ar^n\] the sum is \[\frac{a}{1-r}\]